# Streamlining Redundant Layers to Compress Large Language Models

### v4

@@ Abstract
This paper introduces LLM-Streamline, a pioneer work on layer pruning for large language models (LLMs). It is based on the observation that different layers have varying impacts on hidden states, enabling the identification of less important layers to be pruned. LLM-Streamline comprises two parts: layer pruning, which removes consecutive layers with the lowest importance based on target sparsity, and layer replacement, a novel module that trains a lightweight network to replace the pruned layers to mitigate performance loss. Additionally, a new metric called stability is proposed to address the limitations of the widely used accuracy metric in evaluating model compression. Experiments show that LLM-Streamline outperforms both previous and concurrent state-of-the-art pruning methods in terms of both performance and training efficiency.
@@ 4.1 Setup
To validate the effectiveness of our proposed method, we conduct experiments on popular open-source LLMs, including Llama2-7B and Llama2-13B  (Touvron et al., 2023). Following previous work (Men et al., 2024; Yang et al., 2024), we perform experiments with a 25% compression rate and extract data from SlimPajama for layer pruning and layer replacement, where SlimPajama is a dataset containing data from different domains for pre-training LLM. Sheared LLaMa  (Xia et al., 2023) finds that the performance degradation of pruned models varies across different domains, and proposes determining the distribution of data from different domains based on the degree of performance degradation. Therefore, we randomly sample the data based on the distribution used by Sheared LLaMa  (Xia et al., 2023), finally constructing the dataset containing 30,000 pieces of data. We randomly select 500 samples from this dataset and input them into LLMs, generating Fig. 2, and use these 500 data samples for layer pruning. All 30,000 pieces of data are used to train the lightweight network. We utilize two types of lightweight networks: a Feed-Forward Neural Network (FFN), referred to as Ours (FFN), and a Transformer Layer, referred to as Ours (Layer). The FFN is randomly initialized, while the Transformer Layer inherits the parameters from the first pruned layer. Additionally, we explore a purely pruning approach without incorporating any lightweight network, denoted as Ours (None). Further experimental details are available in the Appendix C.1.
@@ 4.2 Benchmark
To comprehensively assess the changes in LLM capabilities before and after pruning, we use 12 natural language understanding and question-answering benchmarks for evaluation: CMNLI  (Xu et al., 2020), HellaSwag(HeSw)  (Zellers et al., 2019), PIQA  (Bisk et al., 2020),CHID  (Zheng et al., 2019), WSC  (Levesque et al., 2012),CommonSenseQA(CoQA)  (Talmor et al., 2018), BoolQ  (Clark et al., 2019),MMLU  (Hendrycks et al., 2020), CMMLU  (Li et al., 2023),Race-High/Middle  (Lai et al., 2017), C3  (Sun et al., 2020). The tasks in these benchmarks are formalized as classification tasks, so we refer to these benchmarks as classification benchmarks. For these benchmarks, we use both accuracy and stability as metrics for evaluating the models. Additionally, we include 3 benchmarks: XSum  (Narayan et al., 2018), GSM8K  (Cobbe et al., 2021) and StrategyQA  (Geva et al., 2021), to demonstrate the LLM’s performance on generation tasks after pruning. We refer to these tasks as generation benchmarks. Following the evaluation framework of OpenCompass (Contributors, 2023), we use accuracy to evaluate StrategyQA and GSM8K, and use ROUGE1 to evaluate Xsum.
@@ 4.3 Baseline
4.4 Main Results
We present accuracy and stability for different methods on the classification benchmarks in Table 2 and Table 3, respectively, and Table 4 for the generative benchmarks. The results demonstrate that our proposed LLM-Streamline consistently outperforms the baseline methods, effectively retaining the majority of the capabilities of LLMs across various tasks. Specifically, in classification tasks, LLM-Streamline surpasses LLM-Pruner by 7% in accuracy and 12% in stability on Llama2-7B, and by 16% in accuracy and 19% in stability on Llama2-13B. LLM-Streamline also surpasses LaCo by 5% in accuracy on Llama2-7B. For generation tasks, LLM-Streamline retains nearly 77% of Llama2-7B and Llama2-13B’s capabilities, significantly outperforming other pruning methods. We find that almost all of the pruning methods fail on the GSM8K dataset, which reflects the difficulty of math word problems. However, we find that sufficient training can gradually restore the model’s performance on math tasks, and the specific experimental results are shown in Table 21 of Appendix D.5.
Additionally, comparing the average stability (Average) in Table 3 with the retrained performance (RP) in Table 2 across different pruning methods reveals that stability is often much lower than accuracy. We also observe that accuracy on Race-M and Race-H even increases after model pruning. Furthermore, we find that without using any lightweight network, Llama2-13B achieves the highest accuracy on classification benchmarks, but its stability on classification benchmarks and performance on generation benchmarks are lower. These results indicate that pruned models tend to make correct guesses on some classification questions that they are uncertain, highlighting the limitations of accuracy as a sole measure of pruning method performance. We also conduct experiments on OPT-1.3B, OPT-2.7B, OPT-6.7B, Baichuan-7B, Baichuan-13B, Baichuan2-7B and Baichuan2-13B. Details can be found in Appendix D.1, Appendix D.2, Appendix D.3 and Appendix D.4.
4.5 Impact of Different Lightweight Networks
While FFN achieves the best result, Transformer layer still has performance potential. We perform experiments with Llama2-7B using various lightweight network architectures, including Feed-Forward Neural Networks (FFN), SwiGLU-based Feed-Forward Neural Networks (SwiGLU), and Transformer layers. We also explore various initialization methods for the Transformer Layer, including random initialization (Layer-Random), inheritance of the first pruned layer (Layer-First), inheritance of the last pruned layer (Layer-Last), and averaging the pruned layers (Layer-Avg). The average accuracy and stability metrics across all the classification benchmarks are presented in Table 5, with detailed results on each benchmark available in Appendix D.6. The results show that FFN achieves the best results. Meanwhile, for the Transformer Layer, inheriting the pruned first layer yields the best results. In contrast, the performance of Layer-Avg, inspired by LaCo, shows that averaging weights does not achieve the same effectiveness as the pruned first layer.
In addition, we plot the validation loss curves during the training process for different lightweight networks, as shown in Fig. 3. We can observe that FFN and SwiGLU have already converged by the 10th epoch, whereas the loss of Transformer Layer is still decreasing. This indicates that the Transformer layer still has potential, and further training could yield better results, but this would require more computing resources.
4.6 Impact of Different Pruning Ratios
The performance of the pruned model is linearly correlated with the number of parameters. To verify the model’s performance at different pruning ratios, we evaluate our method not only at the approximately 25% pruning ratio but also at ratios of around 2%, 11%, and 20% on Llama2-7B. The average stability and accuracy metrics across all the classification benchmarks are shown in Fig. 4, with details on each benchmark presented in Appendix D.7. By comparing the performance of the original Llama2-7B, TinyLlama-1.1B, OpenLlama-3B-v2, and Llama2-7B pruned at various ratios, we observe a linear correlation between the performance of both the pruned models and the pre-trained original models relative to the number of parameters. This suggests that the performance of models pruned using our method is comparable to that of pre-trained models with the same number of parameters.
4.7 Comparison of Layer Replacement and LoRA
Layer Replacement outperforms LoRA in both performance and GPU memory consumption. We compare the performance of layer replacement with LoRA. Since layer replacement is trained based on hidden states with a different training objective than LoRA, we additionally train one epoch using the language model loss for layer replacement when comparing it with LoRA. The training details can be found in the Appendix C.2. For layer replacement, we freeze the original model’s weights and train only the lightweight network. In the case of LoRA, we set the rank to 128 to align the number of parameters trained with those of the lightweight networks. We randomly extract 30,000 samples from SlimPajama-6B for layer replacement training and also test with the entire dataset to evaluate the limited impact of extensive data on performance (details in Appendix D.5). For LoRA, we use 300,000 samples from SlimPajama-6B. Table 6 presents the average accuracy and stability across all classification benchmarks, with detailed results available in Appendix D.8. The findings indicate that layer replacement surpasses LoRA in both accuracy and stability, while also requiring significantly less GPU memory and training data.
Figure 4: (a) Stability of the pruned Llama2-7B at pruning ratios of approximately 2%, 11%, 20%, and 25%. Stability is calculated only for the compressed model compared to the original, so it cannot be computed for TinyLlama-1.1B and OpenLlama-3B-v2. (b) Accuracy of the pruned Llama2-7B at different pruning ratios, compared to the original Llama2-7B, OpenLlama-3B-v2, and TinyLlama-1.1B. Metrics are averaged across classification benchmarks.
@@ C.2 Post Training Details
Appendix D Detailed Experimental Results
D.1 Experimental Results of OPT-6.7B
We also conduct experiments on OPT-6.7B. The experimental results are shown in Table 11, Table 12 and Table 13. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
@@ Appendix D Detailed Experimental Results
D.2 Experimental Results of Baichuan-7B and Baichuan-13B
We also conduct experiments on Baichuan-7B and Baichuan-13B. The experimental results are shown in Table 14, Table 15 and Table 16. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
@@ D.2 Experimental Results of Baichuan-7B and Baichuan-13B
D.3 Experimental Results of Baichuan2-7B and Baichuan2-13B
We also conduct experiments on Baichuan2-7B and Baichuan2-13B. The experimental results are shown in Table 17, Table 18 and Table 19. The results indicate that our proposed LLM-Streamline is superior to the concurrent SOTA method, LaCo.
@@ D.3 Experimental Results of Baichuan2-7B and Baichuan2-13B
D.4 Experimental Results of OPT-1.3B and OPT-2.7B
@@
D.5 Results of Sufficient Post-training

### v5

@@ Abstract
This paper introduces LLM-Streamline, a pioneer work on layer pruning for large language models (LLMs). It is based on the observation that different layers have varying impacts on hidden states, enabling the identification of less important layers to be pruned. LLM-Streamline comprises two parts: layer pruning, which removes consecutive layers with the lowest importance based on target sparsity, and layer replacement, a novel module that trains a lightweight network to replace the pruned layers to mitigate performance loss. Additionally, a new metric called stability is proposed to address the limitations of the widely used accuracy metric in evaluating model compression. Experiments show that LLM-Streamline outperforms both previous and concurrent state-of-the-art pruning methods in terms of both performance and training efficiency. Our code is available at this repository.
@@ 4.1 Setup
We conduct experiments on popular open-source LLMs, including Llama2-7B and Llama2-13B (Touvron et al., 2023). Following previous work (Men et al., 2024; Yang et al., 2024), we perform experiments with a 25% pruning rate and extract data from pre-training dataset SlimPajama which contains data from different domains for layer pruning and layer replacement. Sheared LLaMa  (Xia et al., 2023) finds that the performance degradation of pruned models varies across different domains, and proposes determining the distribution of data from different domains based on the degree of performance degradation. Therefore, we randomly sample the data based on the distribution used by Sheared LLaMa  (Xia et al., 2023), finally constructing the dataset containing 30,000 pieces of data. We randomly select 500 samples from this dataset and input them into LLMs, generating Fig. 2, and use these 500 data samples for layer pruning. All 30,000 pieces of data are used to train the lightweight network. We utilize two types of lightweight networks: a Feed-Forward Neural Network (FFN), referred to as Ours (FFN), and a Transformer Layer, referred to as Ours (Layer). The FFN is randomly initialized, while the Transformer Layer inherits the parameters from the first pruned layer. Additionally, we explore a purely pruning approach without incorporating any lightweight network, denoted as Ours (None). Further experimental details are available in the Appendix C.1.
@@ 4.2 Benchmark
We use 12 natural language understanding benchmarks for evaluation: CMNLI  (Xu et al., 2020), HellaSwag(HeSw)  (Zellers et al., 2019), PIQA  (Bisk et al., 2020),CHID  (Zheng et al., 2019), WSC  (Levesque et al., 2012),CommonSenseQA(CoQA)  (Talmor et al., 2018), BoolQ  (Clark et al., 2019),MMLU  (Hendrycks et al., 2020), CMMLU  (Li et al., 2023),Race-High/Middle  (Lai et al., 2017), C3  (Sun et al., 2020). The tasks in these benchmarks are formalized as classification tasks, so we refer to these benchmarks as classification benchmarks. For these benchmarks, we use both accuracy and stability as metrics for evaluating the models. Additionally, we include 3 benchmarks: XSum  (Narayan et al., 2018), GSM8K  (Cobbe et al., 2021) and StrategyQA  (Geva et al., 2021), to demonstrate the LLM’s performance on generation tasks after pruning. We refer to these tasks as generation benchmarks. Following the evaluation framework of OpenCompass (Contributors, 2023), we use accuracy to evaluate StrategyQA and GSM8K, and use ROUGE1 to evaluate Xsum.
@@ 4.3 Baseline
4.4 Main Results
We present accuracy and stability for different methods on the classification benchmarks in Table 2 and Table 3, respectively, and Table 4 for the generative benchmarks. The results demonstrate that our proposed LLM-Streamline consistently outperforms the baseline methods. Specifically, in classification tasks, LLM-Streamline surpasses LLM-Pruner by 7% in accuracy and 12% in stability on Llama2-7B, and by 16% in accuracy and 19% in stability on Llama2-13B. LLM-Streamline also surpasses LaCo by 5% in accuracy on Llama2-7B. For generation tasks, LLM-Streamline retains nearly 77% of Llama2-7B and Llama2-13B’s capabilities, significantly outperforming other pruning methods. We find that almost all of the pruning methods fail on the GSM8K dataset. However, sufficient training can gradually restore the model’s performance on math tasks, and the specific experimental results are shown in Table 28 of Appendix E.8.
Additionally, comparing the average stability (Average) in Table 3 with the retrained performance (RP) in Table 2 reveals that stability is often much lower than accuracy. We also observe that accuracy on Race-M and Race-H even increases after model pruning. Furthermore, we find that without using any lightweight network, Llama2-13B achieves the highest accuracy on classification benchmarks, but its stability on classification benchmarks and performance on generation benchmarks are lower. These results indicate that pruned models tend to make correct guesses on some classification questions that they are uncertain, highlighting the limitations of accuracy as a sole measure of pruning method performance. We also conduct experiments on OPT-1.3B, OPT-2.7B, OPT-6.7B, Baichuan-7B, Baichuan-13B, Baichuan2-7B, Baichuan2-13B (yang2023baichuan), Llama3.1-8B, Llama3.1-70B (dubey2024llama) and Mixtral-8x7B-v0.1 (jiang2024mixtral). Details can be found in Appendix E.1, Appendix E.2, Appendix E.3, Appendix E.4, Appendix E.5 and Appendix E.6. In addition, we compare the performance of LLM-Streamline with other methods at a higher pruning ratio of approximately 50%, and the results can be found in Appendix E.7.
4.5 Impact of Different Lightweight Networks
While FFN achieves the best result, Transformer layer still has performance potential. We perform experiments with Llama2-7B using various lightweight network, including Feed-Forward Neural Networks (FFN), SwiGLU-based Feed-Forward Neural Networks (SwiGLU), and Transformer layers. We also explore various initialization methods for the Transformer Layer, including random initialization (Layer-Random), inheritance of the first pruned layer (Layer-First), inheritance of the last pruned layer (Layer-Last), and averaging the pruned layers (Layer-Avg). The average accuracy and stability metrics across all the classification benchmarks are presented in Table 5, with detailed results on each benchmark available in Appendix E.9. The results show that FFN achieves the best results. Meanwhile, for the Transformer Layer, inheriting the pruned first layer yields the best results. In contrast, the performance of Layer-Avg, inspired by LaCo, shows that averaging weights does not achieve the same effectiveness as the pruned first layer. In addition, we plot the validation loss curves for different lightweight networks, as shown in Fig. 3. We can observe that FFN and SwiGLU have already converged by the 10th epoch, whereas the loss of Transformer Layer is still decreasing. This indicates that the Transformer layer still has potential, and further training could yield better results, but this would require more computing resources.
4.6 Impact of Different Pruning Ratios
The performance of the pruned model is linearly correlated with the number of parameters at modest pruning ratios. To verify the model’s performance at various modest pruning ratios, we evaluate our method not only at the approximately 25% pruning ratio but also at ratios of around 2%, 11%, and 20% on Llama2-7B. The average stability and accuracy metrics across all the classification benchmarks are shown in Fig. 4, with details on each benchmark presented in Appendix E.10. By comparing the performance of the original Llama2-7B, TinyLlama-1.1B, OpenLlama-3B-v2, and Llama2-7B pruned at various ratios, we observe a linear correlation between the performance of both the pruned models and the pre-trained original models relative to the number of parameters. This suggests that the performance of models pruned using our method is comparable to that of pre-trained models with the same number of parameters.
4.7 Comparison of Layer Replacement and LoRA
Layer Replacement outperforms LoRA in both performance and GPU memory consumption. We compare the performance of layer replacement with LoRA. Since layer replacement is trained based on hidden states with a different training objective than LoRA, we additionally train one epoch using the language model loss for layer replacement when comparing it with LoRA. The training details can be found in the Appendix C.2. For layer replacement, we freeze the original model’s weights and train only the lightweight network. In the case of LoRA, we set the rank to 128 to align the number of parameters trained with those of the lightweight networks. We randomly extract 30,000 samples from SlimPajama-6B for layer replacement training and also test with the entire dataset to evaluate the limited impact of extensive data on performance (details in Appendix E.8). For LoRA, we use 300,000 samples from SlimPajama-6B. Table 6 presents the average accuracy and stability across all classification benchmarks, with detailed results available in Appendix E.11. The findings indicate that layer replacement surpasses LoRA in both accuracy and stability, while also requiring significantly less GPU memory and training data.
Figure 4: (a) Stability of the pruned Llama2-7B at different pruning ratios. (b) Accuracy of the pruned Llama2-7B at different pruning ratios, compared to the original Llama2-7B, OpenLlama-3B-v2, and TinyLlama-1.1B. Metrics are averaged across classification benchmarks.
@@ C.2 Post Training Details
Appendix D Inference speed comparison
As shown in the Table 11, we evaluate the inference speed of various pruning methods at similar pruning ratio when generating sequences of length 128. The results indicate that the acceleration effect of LLM-Streamline is slightly inferior to that of SliceGPT and LLM-Pruner.
Table 11: The inference speed of models pruned using different methods.
Llama2-7B Dense LLM-Pruner SliceGPT Ours(None) Ours(FFN) Ours(Layer)
Pruning Ratio (%) 0.00 24.8 25.4 24.0 25.0 24.0
Inference Speed (tokens/s) 19.87 25.91 27.20 25.68 25.88 25.68
Appendix E Detailed Experimental Results
E.1 Experimental Results of OPT-6.7B
We conduct experiments on OPT-6.7B. The experimental results are shown in Table 12, Table 13 and Table 14. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
@@ Appendix E Detailed Experimental Results
E.2 Experimental Results of Baichuan-7B and Baichuan-13B
We conduct experiments on Baichuan-7B and Baichuan-13B. The experimental results are shown in Table 15, Table 16 and Table 17. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
@@ E.2 Experimental Results of Baichuan-7B and Baichuan-13B
E.3 Experimental Results of Baichuan2-7B and Baichuan2-13B
We conduct experiments on Baichuan2-7B and Baichuan2-13B. The experimental results are shown in Table 18, Table 19 and Table 20. The results indicate that our proposed LLM-Streamline is superior to the concurrent SOTA method, LaCo.
@@ E.3 Experimental Results of Baichuan2-7B and Baichuan2-13B
E.4 Experimental Results of Llama3.1-8B and Llama3.1-70B
We conduct experiments on Llama3.1-8B and Llama3.1-70B. The experimental results are shown in Table 21 and Table 22. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
Table 21: Accuracy of different pruning methods on classification benchmarks by pruning Llama3.1-8B and Llama3.1-70B.
LLM Method Ratio Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama3.1-8B Dense 0.00% 65.3 33.0 73.8 68.2 36.5 69.8 74.7 81.1 71.6 64.5 66.8 52.5 63.2 100
SliceGPT 23.9% 38.4 32.1 21.3 37.8 38.5 38.0 39.9 58.7 21.9 23.3 25.8 25.2 33.4 52.8
Ours (None) 24.4% 42.3 33.7 19.3 52.3 36.5 30.7 28.4 58.9 36.6 33.3 39.1 34.4 36.9 58.4
Ours (Layer) 24.4% 55.9 34.5 54.5 67.6 36.5 62.5 62.6 74.5 64.8 55.9 64.9 51.5 57.1 90.6
Llama3.1-70B Dense 0.00% 74.8 33.0 81.6 76.5 37.5 73.0 79.9 83.9 86.8 80.5 79.3 68.8 71.3 100
SliceGPT 29.1% 40.4 31.9 18.9 37.8 37.5 41.0 45.3 61.0 24.1 24.8 37.5 30.5 35.9 50.4
Ours (None) 30.3% 66.1 37.5 58.1 69.0 46.2 61.8 68.4 75.7 81.7 73.2 70.4 62.0 64.2 90.0
Ours (Layer) 30.3% 68.9 34.7 70.0 72.5 42.3 68.9 74.4 79.3 86.8 81.5 78.6 68.2 68.8 96.5
Table 22: Stability of different pruning methods on classification benchmarks by pruning Llama3.1-8B and Llama3.1-70B.
LLM Method Ratio Benchmarks Average
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama3.1-8B SliceGPT 23.9% 63.2 37.0 38.5 35.6 98.1 60.7 59.3 71.2 39.0 46.1 42.1 47.4 53.2
Ours (None) 24.4% 59.9 47.0 39.0 57.0 100 49.8 43.4 64.7 53.4 54.5 57.8 61.4 53.7
Ours (Layer) 24.4% 78.5 49.7 59.9 75.0 100 80.3 84.8 86.5 87.3 86.1 90.8 89.1 80.7
Llama3.1-70B SliceGPT 29.1% 55.6 45.4 32.1 36.1 98.1 58.8 59.6 72.0 32.9 39.6 48.7 45.9 49.3
Ours (None) 30.3% 77.6 43.6 64.5 73.0 91.4 76.8 84.6 85.0 92.4 89.3 84.1 81.4 78.6
Ours (Layer) 30.3% 86.7 95.7 76.1 77.7 95.2 89.5 92.6 93.4 97.2 96.3 95.9 94.6 90.9
E.5 Experimental Results of Mixtral-8x7B-v0.1
We conduct experiments on Mixture of Experts(MoE) model Mixtral-8x7B-v0.1. The experimental results are shown in Table 23 and Table 24.
Table 23: Accuracy of LLM-Streamline on classification benchmarks by pruning Mixtral-8x7B-v0.1.
LLM Method Ratio Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Mixtral-8x7B-v0.1 Dense 0.00% 54.1 33.0 48.6 68.4 56.7 68.2 76.2 81.7 72.4 70.9 71.3 52.8 62.9 100.0
Ours (None) 24.9% 39.0 33.0 26.9 62.8 39.4 45.6 55.4 70.2 41.3 43.3 67.7 39.2 47.0 74.7
Ours (Layer) 24.9% 51.2 34.4 41.5 66.3 56.7 62.0 68.3 77.9 54.5 55.7 69.9 50.2 57.4 91.3
Table 24: Stability of LLM-Streamline on classification benchmarks by pruning Mixtral-8x7B-v0.1.
LLM Method Ratio Benchmarks Average
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Mixtral-8x7B-v0.1 Ours (None) 24.9% 66.9 98.9 62.4 76.4 38.5 69.3 72.4 80.6 57.0 60.3 79.1 64.4 68.9
Ours (Layer) 24.9% 83.6 87.6 76.2 81.3 100 86.8 90.4 90.8 67.8 69.0 85.3 80.4 83.3
E.6 Experimental Results of OPT-1.3B and OPT-2.7B
@@
E.7 Experimental Results of Llama2-7B at around 50% pruning Ratio
We conduct experiments on Llama2-7B at a higher pruning ratio of approximately 50%. The experimental results are shown in Table 26 and Table 27. The results indicate that our proposed LLM-Streamline is superior to the previous SOTA method.
Table 26: Accuracy of different pruning methods on classification benchmarks by pruning Llama2-7B at a higher pruning ratio of approximately 50%.
LLM Method Ratio Benchmarks Average RP
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
LLMPruner 49.2% 26.4 33.2 17.3 43.5 38.5 25.4 32.7 59.2 22.3 21.7 23.4 24.8 30.7 63.4
SliceGPT 48.3% 26.5 32.1 15.4 38.1 42.3 28.1 30.9 53.6 23.6 23.1 25.2 25.3 30.4 61.8
Ours (None) 48.0% 33.1 34.0 17.3 55.4 36.5 31.0 34.3 56.3 26.8 27.2 34.9 27.9 34.6 70.3
Ours (Layer) 48.0% 39.7 33.0 27.7 62.1 36.5 44.3 45.2 63.6 24.6 24.3 33.2 28.1 38.5 78.3
Table 27: Stability of different pruning methods on classification benchmarks by pruning Llama2-7B at a higher pruning ratio of approximately 50%.
LLM Method Ratio Benchmarks Average
C3 CMNLI CHID BoolQ WSC CoQA HeSW PIQA Race-M Race-H MMLU CMMLU
Llama2-7B LLMPruner 49.2% 63.6 96.2 60.8 33.8 87.5 51.5 55.5 70.9 47.8 48.1 56.9 58.1 60.9
SliceGPT 48.3% 48.0 34.9 54.7 19.0 83.7 57.7 52.7 63.4 71.2 61.5 52.2 58.1 54.8
Ours (None) 48.0% 73.2 63.0 62.2 65.2 95.2 52.2 52.5 62.6 48.2 52.5 61.4 58.8 62.3
Ours (Layer) 48.0% 80.8 100 67.7 79.5 95.2 65.3 66.1 76.0 58.7 57.3 63.8 59.4 72.5
E.8 Results of Sufficient Post-training